OpenStudy (anonymous):

Can someone help me to find the points of these equations and tell me what the domain and range of both? Thank you! 1st Equation: y = │x-1│ + 2 2nd Equation: x= -√y

OpenStudy (mathstudent55):

1. Make a table. Choose a few values of x. Calculate y. Then plot. x | y ----- -4 | -3 | -2 | -1 | 0 | 1 | 2 |

OpenStudy (anonymous):

I did that and my teacher said it was wrong? I wrote the following: for the 1st EQ. x 0 1 2 3 y = │x-1│ + 2 3 2 3 4

OpenStudy (mathstudent55):

x | y ----- -4 | -3 | -2 | -1 | 0 | 3 1 | 2 2 | 3 4 | 4

OpenStudy (mathstudent55):

Those are the points you did using my table. If you find the y values for the other x values I suggested, you'll get a better idea of what the graph will look like.

OpenStudy (anonymous):

No, those are the points that I used on my assignment and I all ready know that they are incorrect.

OpenStudy (anonymous):

So using your table it would look like this: x = 0 1 2 3 y= 3 2 3 4 Correct?

OpenStudy (mathstudent55):

They are correct.

OpenStudy (anonymous):

You are saying that the table you have just written is correct, right!

OpenStudy (mathstudent55):

Yes. Your values are correct. I just think you need a few more values using negative values of x, as I put on my table, to get a better idea of what the graph will be like. Then from the graph you can see what the domain and range are.

OpenStudy (anonymous):

I included my graph with my work! So using your number what would he range and domain be on the first equation?

OpenStudy (mathstudent55):

Copy and paste the table below and fill in the missing y-coordinates. x | y ----- -4 | -3 | -2 | -1 | 0 | 3 1 | 2 2 | 3 4 | 4

OpenStudy (anonymous):

I am all ready confused. I found what I felt were the right points created a graph and included what I thought was the correct range and domain, clearly I was wrong and I don't want to get my second try wrong, can you tell me what the correct coordinates are?

OpenStudy (mathstudent55):

The points you showed me are all correct. Doing it it the same you way you found your points, find the y-coordinates for x values from -4 to -1 and fill in my table.

OpenStudy (anonymous):

If i am correct why did my teacher tell me that I am incorrect? I do not understand.

OpenStudy (mathstudent55):

Your points were correct, but they may not be enough to give a complete picture of what the graph looks like. You need more points. Tat's why I suggested doing those negative values of x too.

OpenStudy (anonymous):

He also said all of my points were wrong!

OpenStudy (anonymous):

I included a 3 complete graphs with the points on them!

OpenStudy (mathstudent55):

I don't know what you told your teacher about domain,range, a graph if any, so I can't commnt o that.

OpenStudy (anonymous):

Domain={0,1,2,3} Range ={2,3,4},

OpenStudy (anonymous):

I wrote 3 once since it was repeated and wrote that in a note.

OpenStudy (mathstudent55):

I see now. That is not correct because you are only taking into account the points you used. This graph goes on forever. If you want me to explain it to you, please do what I asked you to do. Fill in my table. Then we'll graph it and I'll explain to you why your domain and range are not correct, and what the correct domain and range are. First please fill in my table.

OpenStudy (anonymous):

x -4 -3 -2 -1 y = │x-1│ + 2 3 2 3 4

OpenStudy (mathstudent55):

Why do you keep doing it that way with points that don't line up. My table is much easier to read.

OpenStudy (mathstudent55):

(-1, 4) is correct. The others are not.

OpenStudy (anonymous):

The example for this assignment looks like this. I think this is what you are asking me for. x y -4 4 -3 3 -2 2 -1 1

OpenStudy (mathstudent55):

y = |x - 1| + 2, for x = -4 y = |-4 - 1| + 2 y = |-5| + 2 y = 5 + 2 y = 7, so you get (-4, 7)

OpenStudy (mathstudent55):

y = |x - 1| + 2 x | y ----- -4 | 7 -3 | 6 -2 | 5 -1 | 4 0 | 3 1 | 2 2 | 3 4 | 4 These are the correct y values for x = -4 to x = -1

OpenStudy (mathstudent55):

Now we plot these points.

OpenStudy (mathstudent55):

Here's our graph "paper." Now we plot the points.